OpenStudy (anonymous):

Choose the equation of the horizontal line that passes through the point (-8, -7).

OpenStudy (anonymous):

?

OpenStudy (anonymous):

]Identify the correct x intercept and y intercept of the equation x + 7y = -21. x intercept = (3, 0); y intercept = (0, 21) x intercept = (-3, 0); y intercept = (0, -21) x intercept = (21, 0); y intercept = (0, 3) x intercept = (-21, 0); y intercept = (0, -3)

OpenStudy (anonymous):

you just have to put values of x and y and see if they satisfy the equation like for x+7y=-21 for x intercept put x=3 and y =0 and for y intercept put x=0 and y =3

OpenStudy (anonymous):

so whats the answer?

OpenStudy (anonymous):

solve it?